200ml of 0.005M AgNO3 reacts with 300ml of 0.01M KCl. Ksp of AgCl 1.8 X 10exp(-10). What is the maximum concentration of Ag+ in the mixture

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
 
AgNO3 + KCl ---- AgCl + KNO3
Conc. of AgNO3 = 200*0.005 = 1 milli mole
Conc. of KCl = 300*0.01 = 3 milli mole
That gives conc. of Cl= 4 millimole
now as,
(Ag+)(Cl-) = 1.8*10^-10
and so,
(Ag+) = 1.8*10^-10 / 4*10^-3
that gives Ag+ = 4.5*10^-3
 
 
Regards
Arun (askIITians forum expert)